# A voice for minorities: diversity in approval-based committee elections under incomplete or inaccurate information

### v1

@@ Appendix C Additions to Section 5
Appendix D Discussion on experiments (Section 6)
We have pre-processed the Polis data before use. First, of all 20 accessible Polis data sets, we remove the london.youth.policing dataset due to a low number of participants, and the bg2025-volunteers dataset because it contains far more statements than voters. Second, as anticipated, participating voters typically vote on only a small portion of the statements, so we obtain a sparse comment-participant-matrix. Since these discussions have been completed already, we will not be able to query voters adaptively, so, for the sake of running our querying algorithms, we need to artificially complete the data. We set all unanswered voter-candidate pairs to disapprove. Third, we removed statements with more than half of voters approving, as they made attaining a high Max Cover score trivial in some datasets.

### v2

@@ Appendix D D   Inaccurate information: proofs of Theorem 7 and Theorem 8
Appendix E E   Discussion on the experiments
We discuss two aspects of the experiments. First, the pre-processing applied to the approval elections obtained from Polis, and second, how we determined the parameter $\phi$ that we used to instantiate the $(q,\phi)$-resampling model of [33] to generate artificial approval elections.
Pre-processing of the data
We did four pre-processing steps. First, of all 20 accessible Polis datasets, we remove the london.youth.policing dataset due to a low number of participants, and the bg2025-volunteers dataset because it contains far more statements than voters (all other sets have far fewer statements than voters, on average $m/n=0.4$). Second, we removed statements with more than half of voters approving, as they made attaining a high Max Cover score trivial in some datasets (and are generally relatively weak statements). Third, we removed voters that didn’t vote on any statement as well as those who didn’t approve of any statements (as they are unsatisfiable).
Fourth and last, as mentioned in the main text already, we need to complete the data. On average across the 18 datasets, voters were presented with 18.2% of the statements. Their responses were 56.3% approvals, 27.3% disapprovals and 16.4% neutral votes. Since we require every query to be answered either as approval or disapproval we need to fill in unanswered queries and queries answered neutrally. We set all such votes to disapprove, which results in a worst-case for the Max Cover score of the algorithms.
Our motivation for this is twofold. First, we are not convinced the partial ballot is representative of the full ballot, making extrapolation a doubtful strategy. Indeed, in Polis, the query selection is random but weighed, taking into account, among other things, how divisive and recent a statement is, and how often it has answered neutrally. Moreover, often up to a third of voters only respond to a single query (often their own statement), while a small portion of voters vote on hundreds of statements. Second, since the majority of cast votes are approvals, extrapolation will make attaining a high Max Cover score trivial. Taking Max Cover as diversity measurement, as we do, this observation shows a diverse committee is practically achievable in the context of Polis (assuming the partial ballots are representative). But this also means that this setting is not of interest to study empirically. By setting all unknowns to disapprove, we obtain the worst-case Max Cover score across all extensions of the partial ballot (therefore, including the real extension), recovering an interesting empirical case-study: if we achieve good Max Cover score here, we also can in practice. However, note that although this is the worst-case for the Max Cover score of each individual algorithm, it is not necessarily worst-case with respect to the ratio between complete/accurate and incomplete/inaccurate algorithms. This would only be true in case the relative decrease in the incomplete/inaccurate Max Cover score is smaller than that of the complete/accurate Max Cover score.
Determining $\phi$ of the Polis data
To determine $\phi$, we compute the approvalwise vector [33, Definition 3] for each election, and then compute the approvalwise distance [33, Definition 4] between each election and the limit of the approvalwise vector [33, Section 5] of $(q,\phi)$-resampling elections for $q=0.0891$ and $\phi=0,1/100,2/100,\dots,1$. For each election, we determine the value of $\phi$ that minimizes this distance and then average over the 18 Polis datasets. This yields $\phi=0.693$.